Find the acceleration of the three weights shown in fig. Assume pulleys are frictionless. Also, find the tension in the cables. 


1
Expert's answer
2021-05-07T08:12:43-0400

"200-T_1=m_1a_1"

"40-T_2=-m_2(a_1+a_2)"

"60-T_2=m_3(-a_1+a_2)"

"-T_1+2T_2=0"


From this system of linear equations


"a_1=3.51\ (m/s^2)"

"a_2=2.7\ (m/s^2)"

"T_1=129.7\ (N)"

"T_2=64.86\ (N)"